the perimeter of a square is 72. Express the length of a diagonal of the square in simplest radical form.

Answer by stanbon(75887)

the perimeter of a square is 72. Express the length of a diagonal of the square in simplest radical form.
----
Each side is 72/4 = 18
----
Use Pythagoras to get:
diag^2 = 18^2 + 18^2
---
diag^2 = 2*18^2
---
diag = 18*sqrt(2)
========================
Cheers,
Stan H.